OpenStudy (anonymous):

solve the system of liner equations. -9x+6y=-30 -12x+9y=-30

OpenStudy (radar):

1. -9x +6y =- 30 2. -12x + 9y =-30 Now multiply Equation 2 by a -2/3 getting: 8x -6y = 20 Now add in order to eliminate y ----------------- -x = -10 so x = 10 Now substitute value found for x in equation 1. getting: -9(10) + 6y =-30 -90 + 6y = -30 6y = 60 y = 10 How about that both x and y = 10